Assume that you are in charge of setting the price of lawnmowers at the department store you manage. The cost of providing lawnmowers is given by TC = 800 + 70Q, and you estimate that the price elasticity of demand is -2.75. What price should you set in order to maximize profit?
Answer #1

MC = d(TC)/dQ = 70

PE = - 2.75

Price = MR x e/(e - 1)

MR = MC at maximum profit level:

Price = 70 x 2.75/1.75 = 110